### v2

# Multimodal Conversation Structure Understanding

Abstract
Conversations are usually structured by roles—who is speaking, who’s being addressed, and who’s listening—and unfold in threads that break with changes in speaker floor or topical focus. While large language models (LLMs) have shown incredible capabilities in dialogue and reasoning, their ability to understand fine-grained conversational structure, especially in multi-modal, multi-party settings, remains underexplored. To address this gap, we introduce a suite of tasks focused on conversational role attribution (speaker, addressees, side-participants) and conversation threading (utterance linking and clustering), drawing on conversation analysis and sociolinguistics. To support those tasks, we present a human annotated dataset of 4,398 annotations for speakers and reply-to relationship, 5,755 addressees, and 3,142 side-participants.
We evaluate popular audio-visual LLMs and vision–language models on our dataset, and our experimental results suggest that multimodal conversational structure understanding remains challenging. The most performant audio-visual LLM outperforms all vision–language models across all metrics, especially in speaker and addressee recognition. However, its performance drops significantly when conversation participants are anonymized. The number of conversation participants in a clip is the strongest negative predictor of role-attribution performance, while acoustic clarity (measured by pitch and spectral centroid) and detected face coverage yield positive associations. We hope this work lays the groundwork for future evaluation and development of multimodal LLMs that can reason more effectively about conversation structure.
1 Introduction
Multi-party conversation understanding involves identifying and structuring interactions between multiple speakers and recipients in a face-to-face setting. Such conversations typically involve complex floor-claiming and turn-taking patterns marked by various pragmatic cues. Modeling these structures is useful for any dialogue system, and it has the potential to shed light on the power dynamics between interlocutors, reveal implicit social hierarchies, and enhance our understanding of how conversational roles shift over time, among other things. While multimodal large language models (LLMs) have shown potential in reasoning over complex videos and dialogues, whether they can adequately parse the structure of conversation—resolve the reply-to relationship between utterances, or attribute roles like speakers and addressees—remains underexplored. To bridge this gap, in this work we make the following contributions:
Figure 1: Our proposed structured prediction task for multimodal conversation structure understanding. Grounded in sociolinguistic and conversational analysis [1, 2, 3, 4, 5], the task requires predicting, for each utterance in the given clip: the speaker, addressee(s), side-participants, and the utterance it replies to. The example, taken from the Big Bang Theory in TVQA [28, 29], illustrates our unified formulation, which treats conversational role attribution and conversation disentanglement as complementary subtasks for modeling the interactional dynamics of dialogue. Further analysis of this example can be found in §A.1.3.
  • We draw on literature in sociolinguistics [1, 2, 3] and conversational organization [4, 5] to devise a framework for multimodal multi-party conversation structure understanding that seeks to capture the interactional patterns of interlocuters. Building on this framework, we formally introduce a suite of tasks (Fig. 1) focused on conversational role attribution (speaker, addressees, side-participants) and conversation disentanglement (utterance linking and clustering).
  • To support those tasks, we build on TVQA [28, 29] and present TV-MMPC, a human-annotated dataset of 4,378 annotations for speakers and reply-to relationship between utterances, 5,599 addressees, and 3,412 side-participants, available at https://doi.org/10.7910/DVN/4KUKUL.
  • We evaluate popular vision–language models and audio-visual LLMs on this annotated dataset, and find that there remains room for improvement for modern multimodal LLMs: while Gemini 2.0 Flash outperforms vision–language models we study, its performance degrades when conversation participants are anonymized; its superior performance correlates with visual and acoustic clarity. This points to memorization and cross-modal cues, among others, as promising directions for future work in multimodal dialogue understanding.
2 Related work
Multi-party conversation understanding.
Multi-party conversation understanding (i.e., dialogues with more than two participants) can be summarized as solving “who says what to whom” [6]. This work tends to consider a limited set of modalities, including text- [7, 8, 9] or audio-only [10, 11] approaches. Representative tasks include addressee recognition and conversation disentanglement; both require understanding the relationship between speakers—or more generally, conversation participants—and their utterances. They are designed and built upon text-based data, such as chatroom logs [12, 13]. Works involving audio-visual data focus on speaker activity, identity, and visual co-occurrence [14, 15, 16] and their long-range dependencies [17, 18], mostly without resolving the semantics of conversation structure, until more recently [19]. In this work, we believe that multiple modalities are useful in capturing different aspects of the conversation structure, and a robust understanding of conversational dynamics should mirror human understanding of social interactions.
Video understanding.
Video understanding involves reasoning over an extended temporal context that includes a sequence of images and/or audio signals/ It is fundamental to tasks ranging from movie genre classification to scene graph creation [20, 21, 22, 23, 24, 25, 26]. Such work focuses on multimodal representations of video amd parsing interactions between visible objects, exemplified by visual commonsense reasoning (VCR, 27) and TVQA [28, 29], require models to infer plausible interactions between characters in an image (or a sequence of frames sampled from videos) and reason over their actions. However, they typically do not explicitly resolve the underlying conversation structure and often discards speech data (audio or transcription). Those works converge in recent advances of foundational models: they are natively multimodal [30, 31, 32, 33], general-purpose models that can also reason over videos. In this work, we assess the capabilities of such models to resolve the conversation structure of such long-form videos. Related work that bears most resemblance to this study does consider multiple modalities for solving conversation structure (addressee recognition) but involves a controlled setting [34, 35]. This work represents interactions and settings represented on screen, and we discuss the impact of this in §6.
3 The TV-MMPC dataset
In this section, we describe how we construct our TV-MMPC (Multi-modal Multi-Party Conversation) dataset. We start from TVQA [28], which includes sampled video frames, audio of the original clips (60–90 seconds), and the associated subtitles. For annotation, we randomly sample 50 clips from the following shows in TVQA: The Big Bang Theory, Friends, House M. D., and How I Met Your Mother. To adapt TVQA for our purpose, we take the following steps:
Re-transcribe the audio on the sentence-level.
The original subtitles lack clear sentence boundaries, which makes them unsuitable for annotation aimed at understanding conversational structure. To address this, we transcribe the raw audio using Whisper [36] and align the transcription with the original subtitle data from TVQA.
Infer and standardize speaker labels.
We are interested in conversational participants, which, in the context of TV series, are all characters that appear in the clip. For this reason, we focus on post-processing Some of the subtitles from TVQA contain speaker labels; those labels are usually the first names of the characters in the clip. To facilitate downstream face recognition and annotation, we standardize those labels and map them to actor names by querying TMDb [^1] https://www.themoviedb.org/ for the cast list of each episode. This gives us the mapping between the canonical names of the characters and the actors who play them. With this information, we project TVQA speaker labels where possible (i.e., when the label is present and not unknown for the given sentence) and the algorithm described in [20, p. 24] to align TVQA subtitles and Whisper transcriptions. For a second pass, we infer speaker labels from the sampled frames and audio: we use the most frequently occurring face during the speech in question,  and we recognize the face following the pipeline described in [37], which maps the faces to actor names on IMDb, [^2] https://www.imdb.com/ and the character who dominates the duration of the sentence is treated as the speaker of the sentence. Those speaker labels are verified during the annotation stage; if no face or actor is detected from the automatic pipeline described above, we manually annotate them too.
Annotation guidelines and procedure.
For ease of annotation, we stitch the sampled frames with face captions and use the audio to create a rough approximation of the original clip. Annotators are carried out by co-authors of the paper; we watch the clips on a dedicated annotation interface and annotate conversational roles and reply-to for each utterance. For more details, see Supplementary Information.
Dataset statistics.
See Table 1.
Table 1: Summary statistics for human-annotated data of 200 randomly sampled clips from TVQA.
Per clip Overall
# Unique # Avg. Total # Unique # Total
Speakers 4.18 21.89 156 4,378
Addressees 4.48 28.00 168 5,599
Side Participants 2.90 17.06 121 3,412
4 Conversation structure understanding
On a conceptual level, our goal is to answer the following questions for any video clip: a.) Who gets to start a conversation and hold the floor? b.) Who is being directly addressed? c.) Who listens in on the side? In articulating our proposed framework of multimodal conversation structure understanding, we build on two existing lines of work:
Conversation disentanglement.
Here, we want to resolve the conversation structure and detect the start of individual sub-conversations or conversational threads [38, 39]. We follow the formulation in [40]: Given an utterance of interest (UOI) $u_{t}$, the task is to identify its parent utterance $u_{p}$, which is the utterance that $u_{t}$ directly replies to ($u_{t}\rightarrow u_{p}$). In other words, we find the most relevant preceding utterance in the conversational thread; if the utterance marks the start of a new thread, we say it replies to itself. The thread is the transitive closure of such pairwise links; the main task for conversation disentanglement is, then, the reply-to task.
Conversational role attribution.
We draw on the standard tasks of speaker and addressee recognition [41, 42] and describe a more general task of conversational role attribution. According to [43], conversational roles are observably cast or projected by the speaker $s_{t}$ at time $t$, regardless of what happens at $t+1$ (e.g. addressee might not hear or understand the speaker, and that does not impact our annotation). We revise the taxonomy proposed in [44, 45] (see Table 2) by organizing participant roles along three binary dimensions: whether a participant is addressed by the speaker, ratified as a member of the conversational group, and known to be attending or perceiving the utterance. This reformulation clarifies distinctions among roles and highlights their dependency: addressees are explicitly addressed, ratified as participants, and known to be listening; side-participants are not addressed but still ratified and perceptually engaged. Bystanders, by contrast, are neither addressed nor ratified, and may not be known to be present, which, unlike in [45], includes non-active participants who are overhearers.
Table 2: Participant role matrix.
Role addressed ratified known
Addressee $+$ $+$ $+$
Side-participants -- $+$ $+$
Bystanders -- -- $\pm$
Formally, given an input video clip $V$, we first extract the sequence of utterances $\mathcal{U}=\{u_{1},u_{2},\dots,u_{|\mathcal{U}|}\}$, where each utterance $u_{i}\in\mathcal{U}$ corresponds to a segment of the video defined by a start time $t_{s,i}$ and an end time $t_{e,i}$, $t_{e,i}\leq t_{s,i+1}$. We also identify the set of participants $\mathcal{P}=\{p_{1},p_{2},\dots,p_{|\mathcal{P}|}\}$ appearing in the video, which are derived from face recognition or other identity tracking methods applied to $V$. For each utterance of interest (UOI) $u_{i}\in\mathcal{U}$, the goal is to predict its conversational context by identifying its speaker, addressees, side-participants, and the utterance it directly replies-to. A multi-party conversation structure solver $\mathcal{G}$ maps the current utterance and relevant context to these structural elements:
$\mathcal{G}:(u_{i},\mathcal{U},\mathcal{P})\mapsto\{u_{p},S_{i},\mathcal{A}_{i},\mathcal{E}_{i}\}$ (1)
where reply-to utterance $u_{p}\in\{u_{1},\dots,u_{i-1}\}$ is the parent utterance that $u_{i}$ directly replies to, identified by a reply-to function $R:u_{i}\mapsto u_{p}$ such that $p<i$. If $u_{i}$ initiates a new thread, $u_{p}=u_{i}$; speaker $S_{i}\in\mathcal{P}$ is the participant who produced utterance $u_{t}$; addressees $\mathcal{A}_{i}\subseteq\mathcal{P}$ is the set of participants primarily addressed by $S_{i}$ during $u_{i}$; side-participants $\mathcal{E}_{i}\subseteq\mathcal{P}$ is the set of participants actively listening or engaged in the conversation involving $u_{i}$, but not directly addressed by $S_{i}$. Note that $\mathcal{A}_{t}\cap\mathcal{E}_{t}=\varnothing$. Given this formulation, the bystanders are participants who are neither addressees nor side-participants but belonging to the same thread and can therefore be determined heuristically.
5 Experiments
In this section, we evaluate six popular audio-visual models and vision–language models on TV-MMPC (§5.3) and explore resource-efficient approaches for task adaptation (§5.4).
5.1 Evaluation metrics
Following existing work, we use the following metrics to evaluate the performance of models:
5.1.1 Conversational roles
Accuracy.
[…]
Set-Based F1.
Existing work uses F1 scores for addressee recognition, treating it as a multi-class classification task. Our setup allows multiple addressees and side-participants, so we adapt multi-label metrics that compare gold (annotated) and predicted sets of participants: $\nicefrac{{1}}{{N}}\sum_{i=1}^{N}\text{F1}_{i}$, where $\text{Precision}_{i}=\nicefrac{{|\hat{Y}_{i}\cap Y_{i}|}}{{|\hat{Y}_{i}|}},\text{Recall}_{i}=\nicefrac{{|\hat{Y}_{i}\cap Y_{i}|}}{{|Y_{i}|}},\text{F1}_{i}=\nicefrac{{2\cdot\text{Precision}_{i}\cdot\text{Recall}_{i}}}{{\text{Precision}_{i}+\text{Recall}_{i}}}$, and $Y_{i}$ is the gold set and $\hat{Y}_{i}$ is the predicted set for utterance $i$.
5.1.2 Conversational thread
Reply-to involves first identifying the pairwise links between an utterance of interest and its parent utterance (child utterance replies to the parent), and secondly turning those links into conversational threads. We can see the former as equivalent to binary classification (i.e. whether two utterances exhibit the reply-to relationship) and the second as a clustering problem. The former is evaluated using the F1 score; the latter on the cluster-based metrics below:
Normalized Variation of Information (NVI).
This metric measures the dissimilarity between clustering assignments, based on entropy and mutual information:
$\text{VI}(C,C^{\prime})=H(C)+H(C^{\prime})-2I(C,C^{\prime}),\quad\text{NVI}(C,C^{\prime})=1-\frac{\text{VI}(C,C^{\prime})}{\log_{2}N}$ (2)
where $H(C)$ is the entropy of clustering $C$, $I(C,C^{\prime})$ is the mutual information, and $N$ is the total number of utterances. We report $100\times\left(1-\text{NVI}\right)$, so larger is better.
One-to-One Overlap (1–1, [47]).
This metric finds a one-to-one mapping between gold and predicted clusters that maximizes their overlap. It treats the contingency matrix $M$, where $M_{i,j}$ is the number of shared utterances between gold cluster $i$ and predicted cluster $j$, as a weighted bipartite graph. The optimal assignment is solved using a min-cost flow algorithm, which identifies the pairing of clusters that yields the greatest total intersection:
[…]
where $\pi$ is a one-to-one mapping from gold to predicted clusters.
Exact Match (EM, [48]).
This metric computes precision, recall, and F1-score based on exact identity between individual gold-standard clusters $c_{i}\in C_{\text{gold}}$ and predicted clusters $c^{\prime}_{j}\in C_{\text{pred}}$. A match is identified for a gold cluster $c_{i}$ if there exists any predicted cluster $c^{\prime}_{j}$ such that $c_{i}=c^{\prime}_{j}$ (i.e., they contain the identical set of elements). Letting $N_{\text{match}}$ be the total number of unique gold clusters $c_{i}$ for which such a match exists, Precision is defined as $P=N_{\text{match}}/|C_{\text{pred}}|$, Recall as $R=N_{\text{match}}/|C_{\text{gold}}|$, and F1 is their harmonic mean, measuring the perfect recovery of clusters.
5.2 Heuristic baseline
We first establish a heuristic baseline. We use Whisper [36] and pyannote [49] to obtain word-level timestamps and speaker labels. Since pyannote only provides generic speaker labels, we determine the speaker by aggregating the frequency of each face appearing at the word level, then assigning the speaker role to the face that appears most frequently within each sentence. For each utterance $u_{i}$, we consider faces present in the context window $[i{-}1,i]$. Among these, the most frequently occurring face excluding the speaker is designated as the addressee, while all remaining faces are labeled as side-participants. For the linking task, the previous utterance is treated as the parent utterance (i.e., $u_{i}$ always reples to $u_{i-1}$).
5.3 LLM zero-shot performance
Table 3: LLM zero-shot performance on conversational roles and threads prediction. Metrics are reported with 95% confidence intervals (CIs) from 10,000 bootstrap resamples.
Conversational roles Conversational threads Speaker Addressees Side-part. Linking Clustering Acc. Set F1 Set F1 F1 1–NVI 1–1 EM F1 Heuristic baseline 34.67 [29.13–40.29] 19.49 [16.15–23.07] 36.98 [31.97–42.14] 92.67 [90.61–94.60] 83.34 [79.24–87.41] 76.20 [70.08–82.25] 31.93 [19.52–45.92] Vision–language models (image and text) LLaMA 4 Scout 49.23 [43.09–55.33] 38.47 [32.94–44.15] 42.70 [35.04–50.76] 87.69 [85.07–90.19] 82.60 [78.51–86.63] 76.13 [70.19–81.96] 31.01 [18.80–43.96] GPT-4.1 mini 55.76 [49.04–62.32] 46.28 [40.38–52.21] 53.72 [46.36–61.32] 81.40 [78.42–84.25] 78.61 [74.68–82.50] 75.82 [70.90–80.64] 21.68 [11.89–32.49] o4-mini 53.39 [46.72–60.12] 49.37 [43.69–55.09] 56.88 [48.99–64.71] 84.65 [81.50–87.70] 78.15 [74.13–82.17] 74.98 [69.89–80.02] 24.87 [14.79–35.83] Gemini 2.0 Flash 51.78 [45.45–58.08] 43.40 [37.94–48.93] 53.87 [46.45–61.44] 85.46 [82.67–88.14] 80.49 [76.52–84.43] 77.18 [72.08–82.21] 24.37 [13.65–36.26] Audio-visual LLMs (video with audio, image, and text) Qwen 2.5-Omni 7B 32.24 [27.04–37.58] 23.03 [18.45–27.91] 25.94 [20.22–32.12] 60.36 [56.06–64.57] 66.66 [62.75–70.67] 62.19 [57.51–66.99] 10.03 [4.53–16.57] Reka-Flash 41.48 [36.60–46.49] 35.71 [30.77–40.69] 15.56 [11.99–19.52] 86.67 [84.38–88.86] 81.87 [77.91–85.73] 74.41 [68.40–80.32] 27.00 [15.50–39.50] Gemini 2.0 Flash 78.60 [74.21–82.61] 68.11 [63.29–72.86] 57.68 [51.20–64.18] 89.51 [87.63–91.32] 85.21 [81.73–88.58] 80.33 [75.32–85.15] 34.60 [23.73–46.01]
We evaluate the zero-shot performance of two types of models on our tasks: vision–language models (LLaMA 4 Scout, [^3] https://huggingface.co/meta-llama/Llama-4-Scout-17B-16E GPT-4.1, [^4] https://openai.com/index/gpt-4-1/ o4-mini, [^5] https://openai.com/index/introducing-o3-and-o4-mini/ and Gemini 2.0 Flash [^6] https://cloud.google.com/vertex-ai/generative-ai/docs/models/gemini/2-0-flash) and audio-visual LLMs (Reka-Flash [32], Qwen 2.5 Omini [50], and Gemini 2.0 Flash with reconstructed clips). The multimodal LLM $\mathcal{F}$ performs a structured prediction $o_{i}$ for each $u_{i}\in\mathcal{U}$:
$o_{i}=(o_{s_{i}},\;o_{a_{i}},\;o_{e_{i}},\;o_{r_{i}})\sim\mathcal{F}(C_{i},\mathcal{P}_{V_{i}},Q_{u_{i}})$ (4)
where $C_{i}$ is the context of $u_{i}$, $\mathcal{P}_{V_{i}}=\{p_{1},p_{2},\dots,p_{|\mathcal{P}|}\}$ is the set of identified participants associated with $C_{i}$, which corresponds to the cast list of the episode from which the clip is taken. Participants include the text of character names and, where available, the best face crop between $t_{s,i}$ and $t_{e,i}$. $Q_{u_{i}}$ is the question associated with the $u_{i}$ in: it includes the line index $i$, the start and end timestamps $t_{s,i}$ and $t_{e,i}$, and the question text itself, which prompts the model for conversation structure analysis. The output ${o}_{i}$ contains the index of the reply-to utterance $o_{r_{i}}\leq i$, identities of the predicted speaker $s_{i}$, addressees $a_{i}$, side-participants $e_{i}$; note $o_{a_{i}},o_{e_{i}}\subseteq\mathcal{P}_{V_{i}}$. Given the design of models, we need to structure the overall prompt differently for each class of models, particularly what context information $C_{i}$ to put in the model:
Vision–language models.
The standard setup in video understanding often involves sampling visual frames from a video while discarding its audio. Here, context $C_{i}=\mathcal{U}^{\star}_{i}$, and $\mathcal{U}_{i}=\{(v^{\star}_{j},u_{j},\;f_{j})\}_{j=1}^{|\mathcal{U}|}$: $v^{\star}_{i}$ is the final frame sampled from $V_{i}$ in the interval $[t_{s,j}\,,t_{e,j}]$, interleaving with the words of $u_{j}$; this is further augmented with audio-derived speaker features $f$ extracted using Librosa [51]: speech rate (slower or faster), pitch range (wide or narrow), pitch mean (lower or higher), and spectral centroid (darker or brighter), which collectively aim to capture the acoustic profile of the speaker.
Audio-visual LLMs.
In this setup, $C_{i}=(V_{i},\mathcal{U}_{i})$, where $V_{i}$ is the input video clip containing utterance $u_{i}$, i.e., its audio-visual context. It is the same clip that annotators watch (described in §3). The transcription $\mathcal{U}_{i}$ consists of utterances spoken in $V_{i}$ ( $u_{i}\in\mathcal{U}_{i}$), each represented as a sequence of words with associated start and end timestamps: $\mathcal{U}_{i}=\left\{\left(u_{j},\;t_{s,j},\;t_{e,j}\right)\right\}_{j=1}^{|\mathcal{U}|}$, where each element corresponds to $j$-th utterance in the clip, and the utterance-level start and end times $t_{s,j}$ and $t_{e,j}$.
The results are reported in Table 3. All vision–language models we consider perform significantly better than the heuristic baseline across all metrics, other than on the reply-to task. Another exception is LLaMA 4 Scout, which falls short on side-participant attribution, although it achieves better results on utterance linking than other models, with a statistically significant margin over GPT-4.1 mini. Among the audio-visual language models we consider, Gemini 2.0 Flash outperforms others by a considerable margin; Qwen 2.5-Omni 7B and Reka-Flash do not show statistically significant improvements over the heuristic baseline on most metrics. These results suggest the need for further research into multimodal models, particularly in developing methods that more effectively leverage and integrate different input modalities.
5.4 Resource-efficient supervised fine-tuning
Table 4: Performance of LoRA SFT of Qwen 2.5-Omni 7B compared to zero-shot. Boldface indicates statistically significant differences, determined by non-overlapping 95% CIs.
[…]
Given the increasing popularity of audio-visual LLMs and the growing number of open-source implementations in this space, we turn to explore task-specific adaptation through supervised fine-tuning (SFT). In this subsection, we describe our approach using Low-Rank Adaptation (LoRA) to fine-tune Qwen 2.5-Omni 7B within the LLaMA Factory framework [52]. To minimize computational overhead, LoRA is applied across all layers of the model with a lightweight rank of 2, and a total training steps of 3,125 on one L40S GPU. Each training sample is limited to 4,096 tokens to manage memory and processing time, and the total number of random samples is capped at 1,000. We employ a leave-one-show-out evaluation strategy: training the model on all but one show, and testing it on the held-out show. This process is repeated for each show in the dataset. The results are reported in Table 3, which we compare the performance before and after LoRA SFT on the same TV series. After fine-tuning, despite the constraint in resources, the model has a statistically significant boost on both addressee and side-participant recognition.
6 Discussion
In this section, we examine key factors that influence model performance of the most capable model from the previous section, Gemini 2.0 Flash. We first analyze the potential impact of participant anonymization and the presence of TV-specific dialogue patterns (§6.1), and then turn to how different modalities of input contribute to or limit the model’s effectiveness across tasks (§6.2).
6.1 Effects of anonymization and non-dialogic lines
Table 5: Impact of anonymization and non-dialogic lines on the performance Gemini 2.0 Flash on conversation structure understanding. Boldface indicates significance by non-overlapping 95% CIs.
[…]
Prior work has shown that anonymizing speakers can influence model learning and downstream performance in text-based dialogue understanding tasks [53, 54]. Motivated by this, we investigate whether anonymizing the identities of conversational participants has a similar impact in our multimodal setting. To do so, for each clip, we create an anonymization mapping based on its individual episode cast list: we assign a random letter to each character within a clip (e.g., “Sheldon Cooper” becomes Character C). We then apply the same frame captioning procedure described in §3, but we replace those actual character names with their anonymized labels. In prompting Gemini, our setup is the same as described in §5.3: the model sees the reconstructed video clips, along with the face crop of the highest confidence score. The difference in performance is presented in Table 5. We observe significant drops in the precision and F1 scores for speaker and addressee recognition, by 64.92 and 52.37 points, respectively.
This points to future work in two directions: first, improving identity-awareness in video understanding [55];  second, assessing the degree to which the performance drop can be attributed to the phenomenon of memorization in the audio-visual LLMs, which is relevant in the context of vision–language models [56] and ASR models [57]. This is more than a straightforward case of test set contamination, where test data overlaps with training data, but a subtler one, where models would leverage their parametric knowledge, among other things, acquired from exposure to cultural artifacts and information about them, potentially across different modalities, when solving a task. Such memorization can impact downstream performance and raises further questions about cultural representation, which has been explored in the context of literary texts [58, 59] and can be relevant to audio-visual LLMs.
Since our dataset is derived from TVQA, during our annotation process, we note that there are edge cases in conversational role attribution that are specific to the medium of television. For example, there are narrators that fill in the background of the story, episode recap segments, or expository monologues for characterization: none of them are part of the actual conversation that takes place in the clip. We annotate those utterances as extra-diegetic (spoken by narrators or characters outside the immediate scene) and monologue (extended speech directed at no one in particular), collectively referred to as non-dialogic lines. While we expect those lines present edge cases and can be challenging to the model, Table 5 (second row) suggests that removing those utterances in evaluation does not lead to significant improvement on those metrics.
6.2 Effects of different modalities of input
Figure 2: Signed explained rank variance from Spearman’s $\rho$ between clip-level features and F1 scores for individual conversational roles. Bars indicate direction and magnitude of correlation: solid ones are significant ($p<0.05$); hatched ones are not. Four categories of features are extracted and presented: those related to length, textual diversity, acoustic clarity, and visual face presence.
To better understand the factors driving performance variation in conversational role attribution, we examine the correlations between interpretable, clip-level features (across textual, acoustic, and visual modalities) and F1 scores (across three roles). We employ Spearman’s $\rho$, a non-parametric measure without linear or normal assumptions, to assess monotonic relationships between ranked variables: a positive $\rho$ indicates that higher values of a feature consistently co-occur with higher ranks of F1 scores, while a negative one suggests higher feature values accompany lower F1 performance.
We present those relationships in Figure 2: for ease of presentation, the impact of each feature is depicted in the signed explained rank-variance: $R^{2}=\text{sign}(\rho)\cdot\rho^{2}\times 100$. Statistically significant correlations ($p<0.05$) are distinguished by bars with solid color fill, while non-significant relationships appear in diagonally hatched bars. It is not surprising that of the features we consider, the number of conversational participants has the strongest negative influence, substantially reducing F1 scores across all roles, which accounts for approximately 3–8% of variance. The influence of clip length is negligible, primarily because TVQA has clips of mostly 60–90 seconds in length.
For textual features, we consider mean sentence length (measured by tokens separated by whitespace) and type–token ratio. We see no significant correlation for this category of features, which suggests that lexical diversity alone does not meaningfully impact model performance. Acoustic features (mean pitch and spectral centroid), indicative of speech clarity and brightness, exert small yet consistent positive correlations with performance, with significance achieved primarily for side-participant attribution (pitch: $\rho=+0.17$, $p=0.018$; spectral centroid: $\rho=+0.15$, $p=0.033$), which suggests that clearer, acoustically brighter speech helps in identifying those peripheral conversational participants. For visual features, we consider the ratio of detected faces relative to the number of unique conversational participants present in annotated data (max face coverage) and raw detected face counts, both capture representational fidelity in the clip. Here we see role-specific dynamics: max face coverage positively correlates with speaker role attribution ( $\rho=+0.15$, $p=0.044$), which sugests accurate face detection can benefit this task. Raw face counts, on the other hand, negatively correlate with side-participant detection ($\rho=-0.23$, $p=0.0012$), which confirms that visual clutter degrades the model performance on this task.
Overall, we observe that features from different modalities contribute differently to role attribution performance: face crowding and acoustic clarity exhibit consistent effects on performance, while textual features offer weaker signals. This suggests that conversation structure understanding is not purely a textual problem: future work on this task might benefit from looking further into grounding in the perceptual context in which conversation unfolds. For future development of audio-visual LLMs, cross-modal cues (e.g. face–participant identity alignment, prosodic salience, etc.) have the potential to enable better parsing of conversation structure.
7 Conclusion
Conversations are usually structured by roles—who is speaking, who’s being addressed, and who’s listening—and unfold in threads that break with changes in speaker floor or topical focus. While large language models (LLMs) are increasingly proficient at dialogue and reasoning, evaluation of their performance on fine-grained conversation structure understanding in multimodal, multi-party settings remains relatively limited. This work is our attempt to bridge this gap and introduce tasks targeting conversational role attribution and conversation disentanglement. To support future work, we present TV-MMPC, a human-annotated dataset built on TVQA, with 4,378 speaker and reply-to annotations, 5,599 adrssee labels, and 3,412 side-participant tags.
We evaluate popular audio-visual LLMs and vision–language models on our dataset, and our experimental results suggest that multimodal conversational structure understanding remains challenging. The most performant audio-visual LLM, Gemini 2.0 Flash, outperforms all vision–language models across all metrics, especially in speaker and addressee recognition. However, its performance drops significantly when conversation participants are anonymized, which points to identity-awareness and memorization in audio-visual LLMs as interesting directions for further research, especially in relation to cultural representation. Additional analysis on Gemini suggests that the number of conversation participants in a clip is the strongest negative predictor of role-attribution performance, while acoustic clarity and detected face coverage yield positive associations. On the other hand, performance of the open-source Qwen 2.5-Omini 7B demonstrates improved performance even with low-rank fine-tuning in a resource-constrained setup, using only a subset of our annotated data. We hope this work lays the groundwork for future evaluation and development of multimodal LLMs that can reason more effectively about conversation structure.
8 Limitations
This work introduces a complex set of tasks that required developing detailed annotation guidelines and extensive annotators, and ensuring quality entails a time-consuming and labor-intensive process. We leave large-scale crowdsourcing, model distillation, or other approaches to scale up annotation, which this work enables along with the annotation guidelines, for future work. To maximize representativeness, we sample clips from TVQA entirely at random. This is not exhaustive of the original dataset and may overlook certain conversational patterns. Relatedly, the TV shows represented in this work is limited to the English language and American culture. We rely extensively on metadata from TMDb and IMDb for actor and character identification; while we believe they are generally reliable and accurate, these sources can include uncredited characters or omit minor roles. The effectiveness of our TVQA processing pipeline is limited by the capabilities of ASR and face recognition models, which is beyond the scope of this work.
Acknowledgments and Disclosure of Funding
The research reported in this article was supported by resources provided by the Google Gemma Academic Program.
Appendix A Supplementary information
A.1 Annotation guidelines
Note to the Reader.
The section outlines the conceptual framework we use to annotate conversational roles and threads in multi-party conversation. [^7] Written primarily by AH and KC, with input and guidance from other co-authors. We draw on such traditions as sociolinguistics, linguistic anthropology, and television studies, especially frameworks developed by Clark and Murphy [44, 45], Sacks and Schegloff [60], and Goodwin [43], who, taken together, lay the foundation of conversation analysis and provide us with the basis of taxonomy on which our role attribution task is based; Goffman [1], on the face-to-face, social nature of dialogic interactions; and finally, McKee [61], who analyzes on TV dialogues in the familiar set of vocabulary (Austinian speech acts, for instance).
Our annotation scheme is the result of several iterations, beginning with a pilot round and subsequent discussions with annotators trained in rhetorics, cognitive science, film studies, and related disciplines. Based on observations from the pilot, we refine the definitions of utterance boundaries, reply-to structure, and role attribution to better reflect how participants—or characters on screen—engage in face-to-face dialogues in TVQA data. We hope the annotation guidelines presented below will bring out the nuances of the annotation tasks at hand and their intellectual stakes, as well as provide the theoretical scaffolding that informs our design choices.
In any dialogic interaction, conversational participants manage and direct attention towards each other, during which they also inhibit various roles (speakers, addressees, etc.). For Goffman [1] in particular, conversations are more than merely linguistic exchanges but ritualized social encounters, regulated by norms that govern such phenomena as turn-taking and floor-claiming. These norms extend beyond linguistic structures alone and are often mediated through non-verbal and non-linguistic cues (gaze, posture, etc.).
If we follow Goffman and attempt to uncover the social architecture of interaction underlying conversational interactions, we can potentially shed light on the tacit rules and alignments that structure how people participate in talk; in the context of media studies, this enables us to probe further into issues pertaining to representation: How do characters include or exclude others? How do they calibrate intimacy, formality, or confrontation? Who gets to speak without contest? Who must listen without reply?
We address three core questions for capturing the dynamics of conversational interactions:
Who initiates the conversation and controls conversational flow (holds the floor)?
This draws on McKee’s notion of speech acts, as characters who initiate dialogue to perform an action, and Goffman’s idea that conversation involves the distribution of attention among its participants. Characters who get to start a conversation and main the floor can be more powerful or otherwise interesting.
Who is explicitly addressed by the speaker?
This reflects Goodwin’s and Clark’s emphasis on speakers and hearers, or Goodwin’s more general inquiry of conversation organization, which helps to shed light on interpersonal relationships and interactive dynamics.
Who participates implicitly as side-participants or bystanders?
Building on Goffman’s notion of adjacency pairs [60] and Clark’s ratified vs. non-ratified participation, participants who are present but not addressed have the potential of illuminating social hierarchies, inclusion/exclusion patterns, and attentional structures within multi-party conversations.
The annotation guidelines are organized as follows: §A.1.1: Definitions; §A.1.2: Annotation procedure; §A.1.3: Examples; §A.1.4: Annotation quality; §A.1.5: Limitations.
A.1.1 Definitions
This is grounded primarily in [1, 61]:
Utterance.
An utterance is a single unit of spoken dialogue, roughly equivalent to a dialogue line. It is a bounded communicative act performed by a speaker within an interaction driven by the need for a plot, of the speaker, or other characters in the context of a TV show. Unlike a statement, which is defined by its propositional content and logical truth conditions, an utterance is defined by its timing, delivery, and interactional role. Utterances need not be syntactically complete sentences, but as moves that accomplish social action. A particular type of utterance is the utterance of interest (UOI), which is the current utterance being annotated. For ease of annotation, we present individual utterances as one sentence transcribed by Whisper [36]. They don’t have to be complete sentences, and they don’t have to be a completed turn in and of itself. This allows us to annotate conversational roles at a reasonable granularity: different lines in the same turn can have different role attributions.
Reply-to.
The reply-to relation connects an utterance of interest to its most relevant preceding utterance, also known as a parent utterance, representing the line to which it responds most directly. This is more prevalent in NLP literature [62], where the task is formalized as a link from a child utterance to its parent, and the relationship is modeled as directed graphs. The default parent utterance is the immediately preceding line; if there is no logical antecedent utterance, the UOI is the beginning of a new thread (more below). Since we present each utterance at the sentence level, a special form of reply-to is that of a continuation: if the speaker is still in the midst of their turn, and the UOI merely continues the previous line, then the previous line is the natural and logical parent utterance. In this view, we might see TV dialogues as linear sequences, where one line necessarily triggers and informs the next, until the current thread runs its course.
Conversational thread.
A coherent sequence of utterances sharing a sustained focus (either on a character or topic), adapted from Goffman’s definition of focused interaction. Dramatic scenes often involve multiple, overlapping threads, each driven by the goals and intentions of characters (aligning with McKee’s concept of dramatic intent). Its annotation involves the following steps:
  • Given each UOI ($u_{t}$), identify the single most relevant preceding utterance as its parent utterance ($u_{p}$).
  • If no suitable preceding utterance exists, annotate the UOI as replying to itself (indicating a new conversational thread).
This annotation captures how dramatic dialogues unfold through pairs of conversational turns, marking shifts in attention, character intention, or conversational control. Those conversational threads (also known as sub-conversations) represent the latent structure of conversations, and the central aim is to segment a broader dialogue into coherent units that represent a stable distribution of attention from conversational participants.
Crucially, our notion of a thread is informed by Goffman’s interactional order, wherein conversation is framed as a ritualized social encounter: participants are not merely exchanging information, but are engaging in a tacit choreography governed by socially shared norms of attention, responsiveness, and turn-taking. These practices include obligations to respond when addressed (more below), to respect shifts in floor control, and monitor others’ contributions for relevance, along with a variety of verbal and non-verbal cues. As such, we emphasize both topical and floor change; indeed, as Goffman notes and demonstrates in his analysis, conversation is not just organized by the semantic content of the utterances. In other words, conversational threading is fundamentally different from, e.g., running topic modeling to cluster utterances: we care about the distribution of attention that we can observe, not that of words. Disentangling a conversation by correctly resolving reply-to relations serves two complementary ends: it, on the one hand, clarifies the information structure of a multi-party exchange; on the other, the observable structure of who replies to whom and when reflects the rules of engagement that make conversation intelligible and socially meaningful.
For more information, see the Appendix of [40], including specific tests and examples.
Conversational roles.
While disentanglement focuses on the structural linkage between utterances, conversational role attribution concerns the dynamic social positions that participants occupy throughout the course of interaction. Each utterance is produced within a shifting ecology of roles, an understanding of who is speaking, to whom, and in front of whom is key to interpreting both literal and, again, social meaning of an interaction.
The first role is the speaker, the animator, the source of the utterance at time $t$. Following Goodwin’s conversational analysis [43] and Clark’s role taxonomy [44, 45], we treat conversational roles as those observably projected at $t$: a speaker’s alignment towards others (signaled through gaze, body orientation, lexical choice) casts other participants into roles. These roles are independent of whether the projected utterance ultimately hears, understands, or replies.
Table 6: Participant role matrix.
Role addressed ratified known
Addressee $+$ $+$ $+$
Side-participants -- $+$ $+$
Bystanders -- -- $\pm$
We annotate each character present in the scene according to three dimensions, which is summarized in Table 6:
  1. 1.
    Addressed: Is the character explicitly spoken to by the speaker?
  2. 2.
    Ratified: Is the character recognized by participants as legitimately part of the conversational group?
  3. 3.
    Perceptually engaged: Is the character known (or shown visually, via multi-modal cues) to be attending to or perceiving the utterance?
This yields the following roles:
  • Speaker: the speaker of the line
  • Addressee: the intended recipient of the line; they are directly addressed, ratified as a co-participant, and likely visibly attending to the speaker, who, in turn, likely constructs the utterance with the addressee in mind, and the addressee is expected to respond or acknowledge the utterance in some form
  • Side-participant: a participant who is ratified, perceptually engaged, but not explicitly addressed. This term is not really used in Goffman, but we prefer it because it has been used in previous work related to ours (e.g., [63]).
  • Bystander: a participant who is neither explicitly addressed nor ratified; may or may not be perceptually engaged (includes overhearers). If someone is not in any of the roles above, they are automatically bystanders, so we are not annotating them specifically.
A.1.2 Annotation procedure
Figure 3: Annotation interface for multimodal conversation structure understanding. Annotators proceed through four steps: a.) watch the reconstructed video clip, b.) verify or correct the speaker label, c.) review and update the reply-to link, and d.) assign addressee(s) and side-participant(s) by dragging the relevant participant badges from the list, derived from the cast list, next to the video.
Annotation will take place on a dedicated interface (Fig. 3), and annotators will follow these concrete steps when examining each utterance in a video clip: based on the list of participants provided to you (derived from the cast list of the episode):
  1. 1.
    Watch the reconstructed clip
  2. 2.
    Identify the speaker and explicitly addressed participants, which usually involves verifying or correcting existing speaker labels
  3. 3.
    Identify the UOI and determine its parent utterance (or mark as new thread, using SELF)
  4. 4.
    Identify participants who are ratified and perceptually engaged but not explicitly addressed (side-participants). Participants are represented as draggable badges, which can be moved from the original list into labeled fields
We include the following tags to handle unidentifiable participants; they are: unknown (not referred to by anyone else, and not identified by our face recognition pipeline, or whose identity is otherwise never revealed to the audience); crowd (a group of unidentifiable people, such as public speaking); OS (off screen and unknown, which presents a special edge case that we might not consider).
In annotating the clips, you are encouraged to re-watch the clip and spend as long as you require, but we recommend a time limit (of, say, 5 minutes) for each clip if this is your first pass so you will not get stuck. Your judgment should be entirely based on the clip and the clip alone.
Figure 4: Annotation enhancements for edge cases and workflow efficiency. Annotators can select characters and special participant tokens for non-standard utterances. Checkboxes allow them to mark lines as continuations (same role as previous) or non-dialogic (extra-diagetic or monologue), with the goal to minimize redundancy and improve overall annotation consistency.
Non-dialogic checkboxes.
Given the narrative form of the TV series, we introduce the following shorthands to indicate utterances that might not appear in typical face-to-face interactions:
  • Extra-diagetic (checkbox): to indicate narrators, “previously on the show”, or other kinds of dialogue lines that are not part of any actual conversation
  • Monologue (checkbox): to better distinguish situations where the character speaks to themselves or other inanimate objects
  • _OS tag: to indicate when the identity of the character is generally known to the audience but not within the scope of the clip.
Those will give us some basis to filter out utterances that are edge cases as we assess the performance of each model on this task. We also have a couple of other life quality features; see Fig. 4.
A.1.3 Examples
For a typical example, let’s consider this scene taken from The Big Bang Theory (season 2, episode 9, “The White Asparagus Triangulation”, segment 02, clip 04). [^8] Segment and clip number are given in TVQA. In it, we observe three characters seated in a row at a movie theater, who are eligible conversational participants: Stephanie Barnett (Sara Rue), on the left, is turned slightly toward the others; Leonard Hofstadter (Johnny Galecki) sits upright, angled more towards Sheldon Cooper (Jim Parsons) on his right:
  1. sheldon
    I’ll find us seats? (#1)                 i
  2. stephanie
    Oh no, we have seats. (#2)                 i
  3. leonard
    Not the right seats. (#3)                 i
Utterance #1 opens the exchange and appears to be phrased tentatively, which suggests a proposal. Its orientation to both Stephanie and Leonard suggests a plural us as the intended group of recipients. We don’t see clearly from the visuals (no direct gaze) or vocative, so we assume both addressees are equally implicated. Being the first utterance in the clip, this is also a thread initiator.
Stephanie provides a reply that rejects Sheldon’s offer (utterance #2); it is triggered and necessitated by utterance #1, which is clearly its parent or reply-to utterance. The “we” still invokes group reference, but the syntactic subject now reclaims authority over the seating decision. Here, we see Stephanie looking directly at, and speaking directly to, Sheldon, which casts Leonard as a side-participant.
Leonard’s utterance (#3) is a reply to #2, seeing as it builds on Stephanie’s claim of having found seats, but disputes its adequacy. Semantically, we can also see this as a response to Sheldon’s original proposal because it supports the need to continue looking for the seats, but structurally, we will annotate the reply-to as #2 for two important reasons: a.) there’s no reason for Leonard to say this if it wasn’t for Stephanie’s utterance #2, which makes the choice of #1 less justifiable as it also overrides the default reply-to; b.) since threads are transitive closures of pairwise relations, and all the utterances respond to Sheldon’s original need to find some seats, those three utterances will form a thread, which signals their connections, and it’s important to distinguish between threads and a single reply-to.
Here are some edge cases:
Incapacitated side-participants.
In this scene from House M. D. (season 4, episode 2, segment 2, clip 11), a team of doctors speak about and to a patient who has locked herself into a room and is experiencing hallucinations. Although she is fully conscious, a named speaking character, and within hearing range of the conversations about her, she is not annotated as a side-participant in those conversations about her symptoms as she is preoccupied by her symptoms, not visually depicted to be listening to the conversation and is not capable of interrupting or joining the ongoing conversation. Later in the scene, however, she is labeled as an addressee after responding to her name.
Off-screen side-participants.
In this scene from House M. D. (season 4, episode 16, segment 2, clip 10), a roomful of doctors speak amongst themselves while one doctor in particular, James Wilson, sits on the side and remains largely unspeaking and unshown. Though the camera is rarely on him and the other doctors address each other rather than him, he qualifies as a side-participant due to being in the same space and scene as the other interlocutors.
Non-human/inanimate objects and monologues.
In this Friends scene (season 10, episode 06, segment 2, clip 17), the speaker addresses a duck, which appears on screen. After discussion, we concluded that addresses must be portrayed by a human in order to be counted and left the addressee blank. The label “monologue” distinguishes between this particular absence of addressee from others which arise in different context (i.e., speaker talking to themselves or having no clear audience).
Extra-diegetic moments.
In this scene from How I Met Your Mother (season 6, episode 9, segment 2, clip 3), speakers cluster around a television to watch a game show, which is then shown on screen with recognizable characters and dialogue. Annotators marked all game show dialogue as extra-diegetic and did not include the TV-watchers as addressees or side-participants.
A.1.4 Annotation quality
We report inter-annotator agreement in Table 7 based on 10–12 sampled clips (not part of the final dataset). Agreement is computed as the average of all pairwise comparisons among the four primary annotators, using the same set of evaluation metrics for experiments, and we observe a clear improvement in the main round of annotation.
Table 7: Annotation quality.
[…]
A.1.5 Limitations
The practical limitations come from our TVQA post-processing pipeline: for instance, annotators observe that cases where conversations are fast-paced (many different speakers, or many different exchanges between two interlocutors, for instance), Whisper would encounter segmentation issues: an utterance would be misattributed to a previous speaker’s line, rather than identified accordingly as an independent utterance (as observed in House M. D., season 5, episode 18, segment 2, clip 20). This error, though infrequent, unsurprisingly suggests that that clips featuring sustained, uninterrupted speech from a single speaker, whether in a series of in-scene utterances or as the result of monologic narration overlaid on a series of images, are more likely to be correctly attributed and segmented than those with many voices, even independent of visual face counts. Future studies may focus on the role of interruption, whether verbal, visual or contextual, on multimodal conversational understandings, as well as what constitutes the boundaries of an utterance as it relates to transcription and segmentation.
A.2 Additional experimental details
class ConversationalRoles(BaseModel):
line_index: int
reply_to: int
speaker: str
addressees: list[str]
side_participants: list[str]
class ClipRoles(BaseModel):
clip_roles: list[ConversationalRoles]
Figure 5: Pydantic schema.
We do not change the default temperature of the LLMs that we evaluate. We use the Pydantic schema for structured output (see Fig. 5). Our system instructions for Gemini 2.0 Flash, similar to those used for other models we consider, are in Fig. 6. This represents a sunccinct summary of our annotation guidelines. For our LLaMA 4 Scout experiments, we run the model in full precision on four H100 GPUs served by vLLM [64], which takes around 3 hours to complete. Qwen 2.5-Omni 7B inference is performed with full precision with Flash Attention 2 enabled [65] on four L40S GPUs, which takes around 7 hours. For Qwen 2.5-Omni 7B LoRA SFT, we follow the instructions and configurations provided on https://github.com/hiyouga/LLaMA-Factory/pull/7537.
You are a video analysis assistant. Your task is to analyze the conversations in a video clip and its associated subtitles. For each dialogue line, you will:
* determine what previous line it is replying to
* determine the speaker, addressees, and side-participants
Heres how to determine the reply-to relationship between utterances to resolve conversational threads:
* The reply-to structure gives us information about floor-claiming and topical change within the clip.
* The character is saying this line because they want to respond to that previous line. What previous line is this current line replying to?
* If the speaker of the last line is the same, you can treat it as continuation and put the index of last line as the reply-to.
* If no previous line triggers this line, then write the current line index, indicating the current line replies to itself, which marks the start of a new conversational thread.
Heres how to determine each role:
* **Speaker:** The character who is speaking the line. Infer this from lip movements, body language, and the context of the dialogue. If a character finishes one line and immediately starts another (very short pause), assume its the same speaker, UNLESS theres a clear visual indication of a scene or speaker change (e.g., a camera cut to a different person starting to speak).
* **Addressee(s):** The character(s) the speaker is *directly* addressing. Use these cues:
* **Eye Contact:** The most important cue. Who is the speaker looking at?
* **Body Orientation:** Is the speakers body turned towards a particular person or group?
* **Dialogue Context:** Does the line contain a name, pronoun ("you"), or clearly refer to a specific individual or group? ("Hey, John..." or "You all need to...")
* **Reactions:** If a character reacts immediately and strongly to a line (e.g., nods, responds verbally, shows surprise), they are likely an addressee.
* If the speaker seems to be talking to everyone present, list all characters who appear to be paying attention.
* If the speaker is talking to a crowd of unidentifiable characters, write "crowd".
* If the speaker is talking to themselves, or no one in particular, write "none".
* **Side-Participant(s):** Any character(s) visible in the scene *during the lines timeframe* who are *not* the speaker or addressees. They are present, and their presence is known to other participants. They can potentially join the conversation at any time.
* If it is not possible to confidently determine if someone is a side-participant, write "unknown".
* If there are no side-participants, write "none".
**Input:**
You will receive a list of subtitle entries. Each entry will be a dictionary with the following keys:
* ‘"line_index"‘: (int) The index of the current entry (subtitle line).
* ‘"start_time"‘: (float) The start time of the subtitle line in seconds.
* ‘"end_time"‘: (float) The end time of the subtitle line in seconds.
* ‘"text"‘: (string) The text of the dialogue line.
You will also receive a list of potential participants for you to assign roles from. You must pick from this list.
With all this information, analyze the video segment corresponding to the start_time and end_time of each subtitle entry.
**Output:**
Provide your output in JSON format, mirroring the structure of the input. For *each* subtitle entry, add the following keys:
* ‘"line_index"‘: (int) The line being analyzed.
* ‘"reply_to"‘: (int) The line index that this current line replies to, could be the same as the current line index or any previous line index.
* ‘"speaker"‘: (string) The name of the speaker. If you cannot determine the speaker, use "unknown".
* ‘"addressees"‘: (list of strings) A list of the names of the addressee(s). This can be an empty list (‘[]‘) if there are no direct addressees, or ‘["none"]‘ if the speaker is speaking generally but to no one in particular.
* ‘"side_participants"‘: (list of strings) A list of the names of the side-participant(s). This can be an empty list (‘[]‘), ‘["none"]‘, or ‘["unknown"]‘
Figure 6: System instruction to Gemini 2.0 Flash.